Fermi National Laboratory
 

Operation Summary for September 7, 2011 - September 9, 2011

 

Wednesday September 7

The day shift began with Operations monitoring the Tevatron (TEV) store 9067 (75.2E30) and the Recycler stash (395.7E10), with the Antiproton Source (Pbar) stacking (9.6E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 9067 at 8:22 AM, put the TEV into shot setup, and established store 9068 at 9:31 AM with an average initial luminosity (AIL) of 361.9E30. Operations turned off the Linac beam so that Linac personnel could begin replacing the power amplifier for a Linac RF station (LRF1). Operators put Pbar into standby during the MI RF LCW system work. A Booster RF person reported that they would do maintenance work on the anode power supplies while Linac was down. A Water Works person reported at 10:31 AM that they had repaired the MI RF LCW leak. RF personnel reported at 11:16 AM that they were changing the modulator for a MI RF station (MIRF3). Stacking, MiniBooNE and NuMI beam returned at 1:15 PM.

During the evening shift, Operations ran beam to users and monitored store 9068.

Thurs day September 8

The midnight shift began with Operations monitoring store 9068 (78.53E30) and the stash (295.6E10), with Pbar stacking (9.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 9068 at 3:58 AM, put the TEV into shot setup, and established store 9069 at 5:33 AM with an average initial luminosity (AIL) of 358.08E30.

During the day shift, Operations terminated store 9069 at 11:47 AM. Tevatron experts then started a period of Tev studies (low energy studies, 3x3’s at 150-450 Gev).

During the evening shift, Operations delivered beam to users and low energy Tevatron studies continue.

Friday September 9

The midnight shift began with Tevatron low energy studies, Operations monitoring the stash (167.3E10), with Pbar stacking (104.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations turned off the Main Injector at 7:19 AM. Experts are accessing the Main Injector to take a look at a kicker (R:KPS3A) which is alarming on a fluorinert low level.

The Plans for Friday and the Weekend

The plans for today are to repair R:KPS3A, continue Tevatron low energy studies and deliver beam to users.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/09/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory